# Strong Linearizability without Compare&Swap: The Case of Bags

### v2

@@ 7.1 Strong-Linearizability
For each $1\leq i\leq s$, let $ins_{i}$ be the Insert operation that performed the last write to Items[$a_{i}$] before $tk_{i}$ read it for the last time (which it did before executing its successful t&s()). Hence, $ins_{i}$ in an Insert($x_{i}$) operation. We begin with proving (i). Let $i$ be an index such that $1\leq i\leq s$. $ins\neq ins_{i}$ since insins was allocated the index $m$ while $ins_{i}$ was allocated the index $a_{i}\neq m$. Since $e$ is a step performed by insins and there is a single producer, $ins_{i}$ was completed in $\sigma$. Hence $ins_{i}$ appears in $S$. Since $S$ is a linearization, in the prefix of $S$ that ends with $ins_{i}$, the bag contains $x_{i}$. By Lemma 21, there is no Take operation after $ins_{i}$ in $S$ that takes $x_{i}$. This implies $x_{i}\in B$.
@@
We start with case (1). $tk_{i}$ read $x_{i}\neq x^{*}$ from Items[$a^{*}$] on LABEL:b-bounded:take:read_item. We showed in the proof of (i) that $ins_{i}$ appears in $S$. If $ins_{i}$ appears in $S$ before $ins^{*}$, then by Lemma 18, a Take operation that performed a successful TS[$a^{*}$].t&s() appears in $S$ after $ins_{i}$. But this is impossible: in the end of the above proof of (i), we showed that no such Take operations appear in $S$ after $ins_{i}$. Therefore, $ins^{*}$ must appear before $ins_{i}$ in $S$. By Lemma 18, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$, and we are done with case (1).
For case (2), $ins^{*}\neq ins$ since the first appears in $S$ and the latter does not. By Lemma 18, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$.
We proceed to case (3). When $ins^{*}$ wrote to Items[$a^{*}$], TS[$a^{*}$] = 0 by Lemma 13. We consider several sub-cases.
@@
In all cases—(3a), (3b) and (3c), $tk^{*}$ takes $x^{*}$ from the bag by Lemma 17. Additionally, $tk^{*}$ appears in $S$ after $ins^{*}$: $tk^{*}$ is linearized after $ins^{*}$ wrote to Items[$a^{*}$] because $tk^{*}$ performs its successful TS[$a^{*}$].t&s() after that write. Since $ins^{*}$ is uncoupled, $tk^{*}$ cannot be linearized before the write to InsertDone by $ins^{*}$, which is the linearization point of $ins^{*}$.
@@
We proceed to the second part, showing that if $x$ is removed from Items, it is removed from the bag beforehand. If Items[a] = $x$ since $w_{InsertDone}$ until the end of $\sigma$, we are done. Else, let $w_{\bot}$ be the step after which Items[a] $\neq x$ for the first time since $w_{InsertDone}$. When the producer writes $x$ to Items[a] at $w$, $a$ is in alloc after it was added to alloc on LABEL:b-bounded:insert:add_to_alloc. To perform another write to Items[a] on LABEL:b-bounded:insert:write_item, the producer needs to choose the index $a$ again on LABEL:b-bounded:insert:choose_index, which requires that $a$ $\notin$ alloc. Before that, $a$ must be removed from alloc on LABEL:b-bounded:insert:remove_from_alloc. Hence, after writing $x$ to Items[a] at $w$, the producer next writes to Items[a] on LABEL:b-bounded:insert:clear_item. Namely, $w_{\bot}$ is a write of $\bot$ to Items[a] by the producer. We will show that a Take operation takes $x$ from the bag of the linearization, after $ins^{\prime}$ inserts it to the bag of the linearization, in a step that precedes $w_{\bot}$. Let $r_{TS}$ be the producer’s step preceding $w_{\bot}$, in which it executes LABEL:b-bounded:insert:check_TS. TS[a] = 1 at $r_{TS}$. By Lemma 13, TS[a] = 0 at $w$. So a successful TS[a].t&s() is performed by a Take operation between $w$ and $r_{TS}$. Let $tk$ be the Take operation that performs the first such t&s, in a step $ts$. tktk is linearized at the latest at $r_{TS}$, hence it is linearized before $w_{\bot}$. It takes $x$ from the bag by Lemma 17. It remains to show that $tk$ is linearized after $ins^{\prime}$. $tk$ is linearized after $w$ because $tk$ is linearized after $ts$ which happens after $w$. Since $ins^{\prime}$ is uncoupled, $tk$ cannot be linearized between $w$ and $w_{InsertDone}$. Therefore, $tk$ is linearized after $w_{InsertDone}$, which is the linearization point of $ins^{\prime}$.
Let $C$ be the configuration that appears in $\sigma$ right before the addition of $m$ to alloc by insins on LABEL:b-bounded:insert:add_to_alloc. Let $B_{C}$ be the set of items in the bag at the end of the linearization of $\sigma$’s prefix that ends with $C$. By Lemma 23, $|B_{C}|$ is less than or equal to the number of non-$\bot$ items that are in Items at $C$. By Lemma 10, the locations in Items with non-$\bot$ values at $C$ are a subset of the locations in alloc at the same configuration, and there are less than $b$ locations in alloc according to the check on LABEL:b-bounded:insert:alloc<b. Hence, $|B_{C}|<b$. No Insert operations are linearized in $\sigma$ after $C$, since the producer is executing the uncoupled Insert insins in this interval, which is linearized only at $e$ that comes right after $\sigma$. Therefore, for any prefix of $\sigma$ that extends the prefix ending with $C$, the bag in the end of the linearization of this prefix does not have more items than $B_{C}$, and we are done.
@@
For each $1\leq i\leq s$, let $ins_{i}$ be the Insert operation that performed the last write to Items[$a_{i}$] before $tk_{i}$ read it for the last time (which it did before executing its successful t&s()). Hence, $ins_{i}$ in an Insert($x_{i}$) operation. We begin with proving (i). Let $i$ be an index such that $1\leq i\leq s$. $ins_{i}$ wrote to InsertDone on LABEL:b-bounded:insert:write_InsertDone in $\sigma$ by definition of $tk_{i}$. Thus, $ins_{i}$ is linearized in $\sigma$, namely, it appears in $S$. Since $S$ is a linearization, in the prefix of $S$ that ends with $ins_{i}$, the bag contains $x_{i}$. By Lemma 21, there is no Take operation after $ins_{i}$ in $S$ that takes $x_{i}$. This implies $x_{i}\in B$.
@@
We start with case (1). $tk_{i}$ read $x_{i}\neq x^{*}$ from Items[$a^{*}$] on LABEL:b-bounded:take:read_item. We showed in the proof of (i) that $ins_{i}$ appears in $S$. If $ins_{i}$ appears in $S$ before $ins^{*}$, then by Lemma 18, a Take operation that performed a successful TS[$a^{*}$].t&s() appears in $S$ after $ins_{i}$. But this is impossible: in the end of the above proof of (i), we showed that no such Take operations appear in $S$ after $ins_{i}$. Therefore, $ins^{*}$ must appear before $ins_{i}$ in $S$. By Lemma 18, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$, and we are done with case (1).
For case (2), $ins^{*}\neq ins^{\prime}$ since the first appears in $S$ and the latter does not. By Lemma 18, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$.
We proceed to case (3). When $ins^{*}$ wrote to Items[$a^{*}$], TS[$a^{*}$] = 0 by Lemma 13. We consider several sub-cases.
@@
In all cases—(3a), (3b) and (3c), $tk^{*}$ returns $x^{*}$ Lemma 17. Additionally, $tk^{*}$ appears in $S$ after $ins^{*}$: $tk^{*}$ is linearized after $ins^{*}$ wrote to Items[$a^{*}$] because $tk^{*}$ performs its successful TS[$a^{*}$].t&s() after that write. Since $ins^{*}$ is uncoupled, $tk^{*}$ cannot be linearized before the write to InsertDone by $ins^{*}$, which is the linearization point of $ins^{*}$.
We proceed to prove (iii). Assume $tk^{\prime}$ exists. Let $C$ be the configuration that appears in $\sigma$ right before the addition of $a^{\prime}$ to alloc by $ins^{\prime}$ on LABEL:b-bounded:insert:add_to_alloc. Let $B_{C}$ be the set of items in the bag at the end of the linearization of $\sigma$’s prefix that ends with $C$. By Lemma 23, $|B_{C}|$ is less than or equal to the number of non-$\bot$ items that are in Items at $C$. By Lemma 10, the locations in Items with non-$\bot$ values at $C$ are a subset of the locations in alloc at the same configuration, and there are less than $b$ locations in alloc according to the check on LABEL:b-bounded:insert:alloc<b. Hence, $|B_{C}|<b$. No Insert operations are linearized in $\sigma$ after $C$, since the producer is executing the coupled Insert $ins^{\prime}$ in this interval, which is linearized only at $e$ that comes right after $\sigma$. Therefore, for any prefix of $\sigma$ that extends the prefix ending with $C$, the bag in the end of the linearization of this prefix does not have more items than $B_{C}$, and we are done.

### v3

@@ 7.1 Strong-Linearizability
For each $1\leq i\leq s$, let $ins_{i}$ be the Insert operation that performed the last write to Items[$a_{i}$] before $tk_{i}$ read it for the last time (which it did before executing its successful t&s()). Hence, $ins_{i}$ in an Insert($x_{i}$) operation. We begin with proving (i). Let $i$ be an index such that $1\leq i\leq s$. $ins\neq ins_{i}$ since insins was allocated the index $m$ while $ins_{i}$ was allocated the index $a_{i}\neq m$. Since $e$ is a step performed by insins and there is a single producer, $ins_{i}$ was completed in $\sigma$. Hence $ins_{i}$ appears in $S$. Since $S$ is a linearization, in the prefix of $S$ that ends with $ins_{i}$, the bag contains $x_{i}$. By Lemma 7.12, there is no Take operation after $ins_{i}$ in $S$ that takes $x_{i}$. This implies $x_{i}\in B$.
@@
We start with case (1). $tk_{i}$ read $x_{i}\neq x^{*}$ from Items[$a^{*}$] on LABEL:b-bounded:take:read_item. We showed in the proof of (i) that $ins_{i}$ appears in $S$. If $ins_{i}$ appears in $S$ before $ins^{*}$, then by Lemma 7.9, a Take operation that performed a successful TS[$a^{*}$].t&s() appears in $S$ after $ins_{i}$. But this is impossible: in the end of the above proof of (i), we showed that no such Take operations appear in $S$ after $ins_{i}$. Therefore, $ins^{*}$ must appear before $ins_{i}$ in $S$. By Lemma 7.9, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$, and we are done with case (1).
For case (2), $ins^{*}\neq ins$ since the first appears in $S$ and the latter does not. By Lemma 7.9, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$.
We proceed to case (3). When $ins^{*}$ wrote to Items[$a^{*}$], TS[$a^{*}$] = 0 by Lemma 7.4. We consider several sub-cases.
@@
In all cases—(3a), (3b) and (3c), $tk^{*}$ takes $x^{*}$ from the bag by Lemma 7.8. Additionally, $tk^{*}$ appears in $S$ after $ins^{*}$: $tk^{*}$ is linearized after $ins^{*}$ wrote to Items[$a^{*}$] because $tk^{*}$ performs its successful TS[$a^{*}$].t&s() after that write. Since $ins^{*}$ is uncoupled, $tk^{*}$ cannot be linearized before the write to InsertDone by $ins^{*}$, which is the linearization point of $ins^{*}$.
@@
We proceed to the second part, showing that if $x$ is removed from Items, it is removed from the bag beforehand. If Items[a] = $x$ since $w_{InsertDone}$ until the end of $\sigma$, we are done. Else, let $w_{\bot}$ be the step after which Items[a] $\neq x$ for the first time since $w_{InsertDone}$. When the producer writes $x$ to Items[a] at $w$, $a$ is in alloc after it was added to alloc on LABEL:b-bounded:insert:add_to_alloc. To perform another write to Items[a] on LABEL:b-bounded:insert:write_item, the producer needs to choose the index $a$ again on LABEL:b-bounded:insert:choose_index, which requires that $a$ $\notin$ alloc. Before that, $a$ must be removed from alloc on LABEL:b-bounded:insert:remove_from_alloc. Hence, after writing $x$ to Items[a] at $w$, the producer next writes to Items[a] on LABEL:b-bounded:insert:clear_item. Namely, $w_{\bot}$ is a write of $\bot$ to Items[a] by the producer. We will show that a Take operation takes $x$ from the bag of the linearization, after $ins^{\prime}$ inserts it to the bag of the linearization, in a step that precedes $w_{\bot}$. Let $r_{TS}$ be the producer’s step preceding $w_{\bot}$, in which it executes LABEL:b-bounded:insert:check_TS. TS[a] = 1 at $r_{TS}$. By Lemma 7.4, TS[a] = 0 at $w$. So a successful TS[a].t&s() is performed by a Take operation between $w$ and $r_{TS}$. Let $tk$ be the Take operation that performs the first such t&s, in a step $ts$. tktk is linearized at the latest at $r_{TS}$, hence it is linearized before $w_{\bot}$. It takes $x$ from the bag by Lemma 7.8. It remains to show that $tk$ is linearized after $ins^{\prime}$. $tk$ is linearized after $w$ because $tk$ is linearized after $ts$ which happens after $w$. Since $ins^{\prime}$ is uncoupled, $tk$ cannot be linearized between $w$ and $w_{InsertDone}$. Therefore, $tk$ is linearized after $w_{InsertDone}$, which is the linearization point of $ins^{\prime}$.
Let $C$ be the configuration that appears in $\sigma$ right before the addition of $m$ to alloc by insins on LABEL:b-bounded:insert:add_to_alloc. Let $B_{C}$ be the set of items in the bag at the end of the linearization of $\sigma$’s prefix that ends with $C$. By Lemma 7.14, $|B_{C}|$ is less than or equal to the number of non-$\bot$ items that are in Items at $C$. By Lemma 7.1, the locations in Items with non-$\bot$ values at $C$ are a subset of the locations in alloc at the same configuration, and there are less than $b$ locations in alloc according to the check on LABEL:b-bounded:insert:alloc<b. Hence, $|B_{C}|<b$. No Insert operations are linearized in $\sigma$ after $C$, since the producer is executing the uncoupled Insert insins in this interval, which is linearized only at $e$ that comes right after $\sigma$. Therefore, for any prefix of $\sigma$ that extends the prefix ending with $C$, the bag in the end of the linearization of this prefix does not have more items than $B_{C}$, and we are done.
@@
For each $1\leq i\leq s$, let $ins_{i}$ be the Insert operation that performed the last write to Items[$a_{i}$] before $tk_{i}$ read it for the last time (which it did before executing its successful t&s()). Hence, $ins_{i}$ in an Insert($x_{i}$) operation. We begin with proving (i). Let $i$ be an index such that $1\leq i\leq s$. $ins_{i}$ wrote to InsertDone on LABEL:b-bounded:insert:write_InsertDone in $\sigma$ by definition of $tk_{i}$. Thus, $ins_{i}$ is linearized in $\sigma$, namely, it appears in $S$. Since $S$ is a linearization, in the prefix of $S$ that ends with $ins_{i}$, the bag contains $x_{i}$. By Lemma 7.12, there is no Take operation after $ins_{i}$ in $S$ that takes $x_{i}$. This implies $x_{i}\in B$.
@@
We start with case (1). $tk_{i}$ read $x_{i}\neq x^{*}$ from Items[$a^{*}$] on LABEL:b-bounded:take:read_item. We showed in the proof of (i) that $ins_{i}$ appears in $S$. If $ins_{i}$ appears in $S$ before $ins^{*}$, then by Lemma 7.9, a Take operation that performed a successful TS[$a^{*}$].t&s() appears in $S$ after $ins_{i}$. But this is impossible: in the end of the above proof of (i), we showed that no such Take operations appear in $S$ after $ins_{i}$. Therefore, $ins^{*}$ must appear before $ins_{i}$ in $S$. By Lemma 7.9, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$, and we are done with case (1).
For case (2), $ins^{*}\neq ins^{\prime}$ since the first appears in $S$ and the latter does not. By Lemma 7.9, a Take operation that takes $x^{*}$ appears in $S$ after $ins^{*}$.
We proceed to case (3). When $ins^{*}$ wrote to Items[$a^{*}$], TS[$a^{*}$] = 0 by Lemma 7.4. We consider several sub-cases.
@@
In all cases—(3a), (3b) and (3c), $tk^{*}$ returns $x^{*}$ Lemma 7.8. Additionally, $tk^{*}$ appears in $S$ after $ins^{*}$: $tk^{*}$ is linearized after $ins^{*}$ wrote to Items[$a^{*}$] because $tk^{*}$ performs its successful TS[$a^{*}$].t&s() after that write. Since $ins^{*}$ is uncoupled, $tk^{*}$ cannot be linearized before the write to InsertDone by $ins^{*}$, which is the linearization point of $ins^{*}$.
We proceed to prove (iii). Assume $tk^{\prime}$ exists. Let $C$ be the configuration that appears in $\sigma$ right before the addition of $a^{\prime}$ to alloc by $ins^{\prime}$ on LABEL:b-bounded:insert:add_to_alloc. Let $B_{C}$ be the set of items in the bag at the end of the linearization of $\sigma$’s prefix that ends with $C$. By Lemma 7.14, $|B_{C}|$ is less than or equal to the number of non-$\bot$ items that are in Items at $C$. By Lemma 7.1, the locations in Items with non-$\bot$ values at $C$ are a subset of the locations in alloc at the same configuration, and there are less than $b$ locations in alloc according to the check on LABEL:b-bounded:insert:alloc<b. Hence, $|B_{C}|<b$. No Insert operations are linearized in $\sigma$ after $C$, since the producer is executing the coupled Insert $ins^{\prime}$ in this interval, which is linearized only at $e$ that comes right after $\sigma$. Therefore, for any prefix of $\sigma$ that extends the prefix ending with $C$, the bag in the end of the linearization of this prefix does not have more items than $B_{C}$, and we are done.